A building is being knocked down with a wrecking ball, which is a big metal sphere that swings on a 10-m-long cable. You are (unwisely!) standing directly beneath the point from which the wrecking ball is hung when you notice that the ball has just been released and is swinging directly toward you. How much time do you have to move out of the way?
Answer #1

time period of big metal ball

T = 2 pi sqrt ( L/g)

T = 2 * 3.14 * sqrt ( 10/9.8)

T = 6.344 second

time we have in our hand to move out is just (1/4) of the time period of complete oscillation

t = T/4 = 1.586 second

==========

do comment in case any doubt, will reply for sure.. Goodluck